# Overcoming the Thermal-Noise Limit of Room-Temperature Microwave Measurements by “Cavity Pre-cooling” with a Low-Noise Amplifier —Application to Time-resolved Electron Paramagnetic Resonance

### v1

@@ III Results
Restricting ourselves to work done on bench-tops at room temperature, we here briefly compare the performance achieved by our demonstration of CPC to that achieved by removing thermal photons via stimulated absorption in an absorptively spin-polarised medium resident inside the cavity; in the literature, this latter approach has been dubbed “masar”[1] or “anti-maser”[24]. Using photo-excited pentacene-doped para-terphenyl as the absorber, Wu et al.[1] cooled a TE01δ mode at 1.45 GHz down to around 50 K -though only in bursts. Using NV- diamond, Ng et al.[2] and Day et al.[5] cooled the same type of mode continuously to 188 K at 2.9 GHz and to 70 K at 9.8 GHz, respectively. Given that these temperatures straddle what we here achieve using CPC, one may conclude that, currently, there is no clear winner. Both approaches can assuredly be improved upon through focussed material science and device engineering. Beyond cost and convenience, a salient advantage of CPC is that the disturbance on the microwave cavity imposed by the cooling mechanism can be removed at the flick of a (microwave) switch. Methods based on internal stimulated absorption are, in contrast, beholden to the particular spin dynamics of the cooling medium. One could however use an external masar/anti-maser (operated either CW or quasi-continuously[40]) as the active cold load (ACL) to implement CPC instead of a HEMT-based LNA, with the connection between the resonator and the anti-maser’s coupling port gated by the microwave switch. As with other pulse-based techniques at microwave frequencies, a successful implementation of CPC relies heavily on the availability of a sufficiently fast, low-loss and glitch-free switch.
To conclude, we have demonstrated a method for pre-cooling the modes of a microwave cavity by temporarily over-coupling them to the input of an LNA serving as an active cold load (ACL). Beyond tr-EPR, other pulsed-EPR techniques such as ESEEM, HYSCORE, and PELDOR could be accelerated by CPC, reducing measurement times[18]. Furthermore, our analysis evaluates how an LNA that is (either continuously or intermittently) coupled directly to the cavity (i.e. with no “protective” isolator inserted in-between) can play an active and beneficial role in reducing the number of thermal photons occupying the cavity’s modes (within the LNA’s bandwidth) -as opposed to just being a passive monitor.
The use of cryogenic ACLs, operating at (say) liquid-helium temperature yet providing input noise temperatures and thus mode cooling down into the mK regime, is perfectly conceivable. Unfortunately, due to self-heating within the device’s active channel/layer, the noise temperatures of what are currently the best semiconductor-based LNAs hit, upon cooling, plateaux of a few K [41, 42]. The construction of effective cryo-ACLs to achieve mK-level mode cooling may thus require new materials and/or architectures (and the associated investment in fabrication equipment to realize such) optimized expressly for mode-cooling –as opposed to the hacking/repurposing of commercially available devices, designed to do something else, like what we have done here. At a more general, sociologically level, it is hoped that this paper induces greater cross-fertilization between pockets of wisdom lurking within microwave radiometry[43], activities across the EPR/NMR communities[44], and hardware innovations in quantum computing/detection (particularly those targeting room-temperature operation)[45].
Acknowledgements: The authors would like to thank Wern Ng and in particular Hao Wu, now at the Beijing Institute of Technology (BIT), for helpful discussions. This work was supported by the U.K. Engineering and Physical Sciences Research Council through Grants Nos. EP/K037390/1 and EP/M020398/1. K.C. acknowledges financial support from the Taiwanese Government Scholarship to Study Abroad (GSSA).
Appendix A SUPPLEMENTARY MATERIAL
A.1 EXPERIMENTAL METHOD - DETAILS
The anatomy of the microwave cavity used in our zero-field tr-EPR spectrometer for cavity pre-cooling demonstrations is shown in Fig. 7. This cavity supports a $\textrm{TE}_{01\delta}$ mode at 1.4495 GHz, exhibiting a loaded quality factor of $Q_{\text{loaded}}=82,000$ when critically coupled, corresponding to an unloaded quality factor of $Q_{0}=164,000$. The cavity consists of five main components: (i) a cylindrical brass enclosure, whose internal walls are plated with silver, (ii) a monocrystalline sapphire ring (supplied by J-Crystal Photoelectric Technology, China) exhibiting low dielectric loss, (iii) an EPR sample, namely a 0.1% pentacene-doped para-terphenyl (PC: PTP) crystal, located in the ring’s bore, (iv) an over-coupling loop (coupling factor = 3.8) connected through a single-pole-double-throw (SPDT) RF switch (viz. a Qorvo RFSW1012PCK-411 eval. board) to a low-noise amplifier (= LNA, viz. a Qorvo QPL9547EVB-01 eval. board), functioning as an active cold load, and (v) a critical-coupling loop connected, via a stub tuner, to a second identical LNA (another QPL9547EVB-01), functioning as the front-end pre-amplifier of a tuned homodyne receiver. The sample receives a Q-switched optical pulse at 532 nm, approximately 5.5 ns in duration and 2 mJ in energy, at a 10 Hz repetition rate, from an integrated Nd:YAG pumped Type II BBO OPO (viz. a Litron Aurora II Integra).
Figure 7: The left image presents the 3D CAD assembly of the microwave cavity (without microwave walls), featuring a movable “ceiling” (= plunger), raised and lowered by a rotating a helical screw (to which the ceiling is mechanically attached) seated in a matching helical thread tapped into the cavity’s “roof” along the cavity’s cylindrical axis. The photo to the right of the drawing shows the cavity’s innards with the ceiling and roof removed. The cavity includes two ports: an over-coupling loop (at 12 o’clock in the photo) and a smaller critical-coupling loop (at 9 o’clock) just poking out of the cavity’s internally-silver-plated wall. The right figure displays the microwave cavity in 2D cross-section. Here, the high-Q sapphire ring is located concentrically on top of a cross-linked polystyrene stand (not shown); its on-axis bore hole, 10 mm in diameter, is filled with the EPR sample (a 0.1% PC:PTP crystal).
The cavity’s internal cylindrical sapphire ring measured 40 mm in height with outer and inner diameters of 68 mm and 10 mm, respectively. The optical “c”-axis of the sapphire (a monocrystal) was aligned with the cylindrical axis. By rotating a screw, the cavity’s axial length (i.e. the internal height of the cavity’s ceiling above its “floor”) could be varied from 150 mm down to 95 mm, allowing wide adjustment in the frequency of the cavity’s $\textrm{TE}_{01\delta}$ mode. The sapphire ring’s hollow supporting pillar was made of cross-linked polystyrene, a solid material offering both low dielectric loss and, unlike PTFE or polyethylene, precision-machineability. This pillar was securely seated into a central hole in the cavity’s floor, holding the sapphire ring 16 mm above the floor. The pillar’s bore allowed for the transfer of pump light, up the cavity’s axis, to the EPR sample. The cavity’s silvered internal surfaces were finely buffed with an abrasive pad, removing patina, then solvent-wiped clean, to leave a gleaming finish.
By constructing a suitable model in COMSOL Multiphysics (a finite-element-based PDE solver) [33], the magnetic mode volume $V_{\text{mode}}$ of the microwave cavity’s $T_{01\delta}$ mode was calculated to be 29 cm3. Experimentally, using a vector network analyser (= VNA, viz. an HP-8753C), the cavity’s loaded quality factor ($Q_{\text{loaded}}=82,000$) and the coupling factors of each of its two ports ($\kappa_{\text{over}}=3.8$, $\kappa_{\text{crit}}=1$) were inferred by the method(s) of Kajfez and Hwan [34]. The PC:PTP crystal, used as our tr-EPR test sample, was grown by the vertical Bridgman method; it fitted into the 10-mm bore of the sapphire ring. The pulsed optical beam from a Q-switched OPO was directed into the microwave cavity via a multimode optical fiber, entering through an 8-mm hole in the centre of the cavity’s floor.
The homodyne receiver for our tr-EPR measurements used a setup akin to Wu et al.’s previous work [50], albeit with the microwave cavity (resonator) now connected to an additional port (for cooling) and with additional routing switches and associated synchronization circuitry -see Fig. 2(c) in the main paper and Fig. 8 below. A sync. pulse from the OPO’s Q-switch driver is used to trigger two pulse generators [both Thandar TG105, one triggered by (= “slaved to”) the other] providing appropriate TTL voltage transitions (“edges”) to connect/disconnect the precooler and switch on/off the EPR spectrometer’s interrogating microwave tone. The most critical switch, needed to perform a clean, abrupt disconnect of the precooling LNA, was a Qorvo RFSW101 SPDT switch, chosen for its low insertion loss, high isolation and sufficient speed; this routes the cooling port’s over-coupling loop first to the active cold load (embodied as the input of a Qorvo QPL9547 LNA) for cooling and, subsequently, to an “open” load (reflecting power back into the cavity, effectively neutralizing the port) for the tr-EPR measurement, when the interrogating microwave tone is applied through the cavity’s critically-coupled monitoring port. To achieve sufficient measurement sensitivity, the homodyne receiver employs a cascade of legacy r.f. amplifiers (total gain $\approx 70$ dB). Upon mixing (= homodyning) down to baseband, the transient tr-EPR signal is channeled (via a DC block) into a fast DC amplifier (viz. a Comlinear E103-I-BNC). All components operate under ambient lab conditions.
Figure 8: The experimental rig (here emphasizing its electrical synchronization and light path) encompassing, from right to left: the OPO’s power supply unit (PSU), the OPO laser head itself, the resonator and it surrounding microwave “plumbing", and the tr-EPR spectrometer’s support electronics.
A.2 CONCEPT AND MODEL OF THE CAVITY-PRE-COOLING TECHNIQUE
A.2.1 Number of photons inside a microwave cavity
Obeying Bose-Einstein statistics, the average number of photons, $\bar{n}$, occupying an electromagnetic mode of frequency $f_{\text{mode}}=\omega_{\text{mode}}/(2\pi)$ inside an isolated (zero ports) cavity maintained at a physical temperature of $T_{0}$, is given by $\bar{n}=\left[\exp\left(\frac{hf_{\text{mode}}}{k_{\text{B}}T_{0}}\right)-1\right]^{-1}$, where $k_{\text{B}}$ and $h$ are Boltzmann’s and Planck’s constants, respectively. Semiclassically, for a microwave cavity with a single port coupling it to an external cold load, as depicted schematically in Fig. 9, the number of photons $q$ occupying the mode as a function of time $t$ will obey

### v2

@@ II Active Cold Loads
Note that the minus sign in front of $T_{\text{min}}$ in Eq. 1 can lead to (partial) cancellation between the two terms on the r.h.s. of the same, all depending on the exact magnitude of $R_{\text{n}}$ and the locations of $S_{11}$ and $\Gamma_{\text{opt}}$ in the complex plane. For the particular LNA used in our room-temperature experiments reported here, namely a Qorvo QPL9547, we calculate, from its specification sheetQorvo (2023), that $T_{\text{cold}}~{}=~{}29.1$ K at 1.45 GHz –around 10 times lower than the LNA’s operating temperature.
Alas, complete sets of scattering and noise parameters are rarely published for cryogenic amplifiers, preventing immediate off-paper predictions of their performance when repurposed as cryo-ACLs. One informative exception is the noise calibration, exploiting a cryogenic impedance tuner, of the type of LNA used for the “ALPACA” phased-array radio telescope, as reported in ref. Sheldon et al. (2021). Though only the magnitude of $S_{11}$ (not its angle) is provided, upon reading off from Figs. 12 and 13, ibid., one can frame this LNA’s available input noise temperature to within: 6.2 K $<T_{\text{rev}}<$ 7.3 K at 1.5 GHz, when operating at 20 K. Equivalent characterizations (including complex scattering parameters) of transistor-based, maser, and parametric amplifiers operating at even lower temperatures would enable the rational design of improved cryo-ACLs.
Figure 2: (a) Conceptual mechanism of active pre-cooling (APC): A microwave resonator is physically maintained at a temperature $T_{\text{hot}}$ (in this article, $T_{\text{hot}}$ equals room temperature, denoted as $T_{\text{0}}$) and supports an electromagnetic mode whose photon occupation number is $\bar{n}_{\text{mode}}$. This mode is coupled to a cold load maintained at a lower temperature $T_{\text{cold}}$ with a coupling factor of $\kappa_{\text{over}}$ between the load and the mode. (b) Simplified equivalent electrical circuit: The resonator, here drawn as an LCR tank circuit, is over-coupled (right) through a (SPDT) switch to either a cold load (in the form of the input of an LNA) or an open circuit –depending on the position of the switch. The mode’s field amplitude (as generates a voltage across the tank circuit) can be measured by a second LNA through its own coupling port (left –this port’s coupling is usually adjusted to be at critical coupling when the cooling LNA is switched out) whose output is recorded on a digital oscilloscope.
Finally, it is instructive to compare Eq. 1 against the standard expression for the effective noise temperature of an LNAWait and Engen (1991); Dunleavy (2009); Modelithics + Mini-Circuits (2015) when used, as intended, as an amplifier:
$T_{\text{eff.}}=T_{\text{min}}+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{(|\Gamma_{\text{s}}-|\Gamma_{\text{opt}}|^{2})}{|1+\Gamma_{\text{opt}}|^{2}(1-|\Gamma_{\text{s}}|^{2})},$ (2)
where $\Gamma_{\text{s}}$ is the coefficient of reflection off the source connected to the input of the LNA –in our case the port of a microwave cavity (see Fig. 1), possibly with a matching circuit/stub interposed. This expression will be needed later on for determining the temperature of a microwave mode –see Supplemental Materialsup (2024). $T_{\text{eff.}}$ can be minimized by “noise matching”, i.e. adjusting $\Gamma_{\text{s}}$ to equal $\Gamma_{\text{opt}}$ in the complex plane (viz. on the Smith chart). Note that minimizing $T_{\text{cold}}$ for an active cold load constructed from an LNA is a different task (their respective formulae 1 and 2 above are different).
III Active Pre-Cooling
The method’s essential idea is to first (i) cool the electromagnetic modes of a microwave cavity containing a sample by over-coupling them to an external active cold load (ACL) and then (ii) swiftly reduce this coupling such that the sample can be interrogated in an electromagnetically cold (and thus quiet) environment before the modes warm back up to ambient temperature. Doing so may enable a more sensitive measurement than otherwise possible.
The first, cooling step is depicted in Figs. 1 and 2(a). In the bench-top experiments we report in this article, the physical walls and contents of the cavity, and the ACL attached to it, all lie at room temperature (nominally, 290 K). Note that the ACL is an electrically powered, single-port device. The output from the LNA that realizes the ACL does not (need to) go anywhere; it can be sunk into a matched passive load at ambient temperature or, if the attenuation backwards through the cooling LNA (essentially $|S12|$) is insufficient, into a second ACL –provided by the input of an additional LNA whose output is terminated by an ambient matched load. [We point out parenthetically that the idea of advantageously terminating the output of the “primary” cooling LNA with a “secondary” cooling LNA (and so on in cascade, as far as needs be) is not newLardizabal (1997).] In our experiments, the input of the cooling LNA is connected via a low-insertion-loss switch to a large loop of wire located inside the cavity, which serves as an over-coupled port.
Figure 3: The cavity’s dissipators, their temperatures, and the resultant noise power outputted.
For understanding and quantifying the effects of coupling an external ACL to a microwave cavity, we now appeal to what Siegman describes simply as a “useful noise theorem” in section 8.3 of his bookSiegman (1964). Consider a cavity whose internal electromagnetic field is monitored by a small (weakly coupled) probe attached to an external port, so providing an output (“O/P”; see Fig. 3 –the equivalent of Fig. 8-10, ibid.). Here, one should regard the cavity’s field as the common “bus” (or “loss-less network” –in Siegman’s more circuit-oriented parlance) through which all of it components, including external loads, interact. Applying the theorem: The available noise power per unit bandwidth $B$ of an electromagnetic mode inside the cavity expressed in units of temperature, namely $T_{\text{mode}}=P_{\text{O/P}}/[k_{\text{B}}B(1-|\Gamma_{\text{O/P}}|^{2})]$, will equal the weighted sum of the temperatures $T_{i}$ of the $N$ bodies that dissipate the mode’s energy, where the weighting $p_{i}$ given to each body equals the fraction of the mode’s energy that it dissipates, and where the sum of these fractions adds up to unity: $\sum_{i}^{N}p_{i}=1$. Mathematically (cf. equation 8-3-3 in Ref. Siegman (1964)): $T_{\text{mode}}=\sum_{i}^{N}p_{i}T_{i}$. We stress that $T_{\text{mode}}$ quantifies the available noise power (per unit bandwidth), taking into account, through the above factor of $(1-|\Gamma_{\text{O/P}}|^{2})]$, the rate at which noise energy is extracted by the probe, where $\Gamma_{\text{O/P}}$ is the reflection coefficient off the probe’s port (again, see Fig. 3). For the sake of the argument here, we assume that the probe’s coupling and thus the noise power the probe itself extracts, as will get sunk into a matched load (at the input of a monitoring instrument), is infinitessimally small, such that this load’s dissipation does not contribute to the $\sum_{i}^{N}$ sum. One must regard the cavity’s enclosing metal walls (of finite conductivity), any dielectric or magnetic material (of finite loss tangent) inside the cavity, plus any other components that contribute to the mode’s intrinsic loss, as such dissipating bodies. External loads coupled to the cavity, such as an active cryogenic load, also count. The thermodynamic temperature that the electromagnetic mode itself adopts is a dissipation-weighted balance between them all.
Over-coupling to an external cold load (passive or active) guaranties that the temperature of a microwave mode is dominated by the load’s available input noise temperature as opposed to the temperature(s) of the cavity’s own intrinsic dissipators. Note that the method affords wide-band cooling: it will cool all of the modes lying within the cold load’s bandwidth, provided they are sufficiently coupled to the load. Quantitatively, what depth of cooling can one expect? Referring back to Figs. 2(̃a) and  3: we suppose for simplicity that all dissipating components of the microwave cavity (grey) are kept at the same common physical temperature of $T_{\text{hot}}=T_{\text{0}}$ (orangey brown). Consider a microwave mode within the cavity that is coupled, with a coupling factor of $\kappa$, to a cold load (blue), whose available temperature is $T_{\text{cold}}$. The dissipation generated in the external load will be $\kappa$ times that of the mode’s own intrinsic (i.e. internal) dissipation. Applying our implementation of Siegman’s useful noise theorem: on average, $\bar{n}_{\text{mode}}$ thermal photons will occupy the mode, corresponding to a mode temperature of $T_{\text{mode}}=hf_{\text{mode}}/[k_{\rm B}\textrm{ln}(1+1/\bar{n}_{\text{mode}})]$, where
$T_{\text{mode}}=\frac{T_{\text{0}}+\kappa T_{\text{cold}}}{1+\kappa}$ (3)
One immediately sees from this equation that, for a high coupling factor $\kappa=\kappa_{\text{over}}\gg 1$, corresponding to a substantially over-coupled port, the microwave cavity’s noise temperature will approach that of the cold load, $T_{\text{cold}}$. The dynamics of the cooling process is treated around Eqs. S1 and S2 in the Supplemental Materialsup (2024), whose steady-state solution (S3) agrees with the thermodynamic treatment above. The time scale on which the mode is cooled is set by the mode’s as-loaded photon lifetime, namely $Q_{0}/[2\pi f_{\text{mode}}(1+\kappa_{\text{over}})]$.
In the case of an active cold load formed by the (gated) channel of the LNA’s input HEMT, the height of this channel is just a few nm (defining a layer of two-dimensional electron gas: a “2-DEG”) and its lateral dimensions (length and width) are sub-micronsBaustista (2008). The cooling capacity of this atto-liter volume of semiconductor is limited. Operationally, one requires that the density, in frequency space, of the electromagnetic modes to be cooled is sufficiently low such that the noise power provided by the total number of modes within the input HEMT’s operating bandwidth is insufficient to saturate/overload its channel. This means that the cavity enclosing and defining the modes must not be too large (to prevent the spectral density of modes becoming too high) nor too hot, and the coupling factors of its modes to the ACL must not be too high. As a last resort, saturation could be thwarted by inserting a low-loss band-pass filter between the cooling LNA and the cavity to act as a pre-selector, so coupling the ACL to fewer modes. A sense of scale is useful with regards to these concerns, however: The noise power impinging on the LNA’s input, at a coupling factor of $\kappa$, from a cavity mode at absolute temperature $T$, equals $2\pi f_{\text{mode}}\kappa k_{\text{B}}T/Q_{0},$ where $k_{\text{B}}$ is Boltzmann’s constant. For the experimental parameters of the demonstrations reported in this article, this amounts to -121 dBm per mode. The combined power from the several dozen modes lying within our LNA’s frequency range (for the dimensions and dielectric loading of the cavity that we used) is thus still around ten orders of magnitude weaker than our LNA’s specified saturation input power (at -1 dB compression). How the LNA’s noise parameters themselves depend on the input (noise) power at elevated power levels is seldom reported, though this huge ratio suggests that the cooling capacity of an ACL constructed from a solid-state LNA enjoys substantial headroom.
Figure 4: (a) Photograph of the experimental resonator’s interior: a sapphire ring holding a 0.1% Pc:PTP crystal, surrounded by the silver-plated walls of a cylindrical brass enclosure. (b) CAD diagram of the same resonator with a two-dimensional axisymmetric simulation of the magnetic flux (white arrows) and magnetic energy density of the TE01δ microwave modeWu et al. (2020a) that this resonator supports. (c) Signal-path anatomy of the experimental time-resolved EPR spectrometer used to demonstrate active pre-cooling. This circuit incorporates homodyne detection of the transient EPR signal. The firing of the optical-parametric oscillator (OPO) laser system and the closing/opening of the two microwave switches are scheduled via transistor–transistor logic (TTL) control pulses (generated by two cascaded pulse generators –both a Thandar TG105) as explained in Fig. 7.
Practicalities (that complicate matters somewhat):